Smart Toys, currently has no debt, expects an EBIT of $45,000 every year forever. Its cost of equity is 16 percent. The corporate tax rate is 30 percent. The company can borrow at 8 percent.

a. What is the current value of the company? 

b. What will the value of the firm be if the company takes on debt equal to 20 percent of its unlevered value? What if it takes on debt equal to 60 percent of its unlevered value? 

c.  What will the value of the firm be if the company takes on debt equal to 55 percent of its levered value? What if the company takes on debt equal to 40 percent of its levered value?   



1
Expert's answer
2022-02-28T16:07:54-0500

a. "45000-45000\times0.3=31500"


b.Let beta be 2

"CARM=2(1+\frac{20}{16})=4.5"

"CARM=2(1+\frac{60}{16})=9.5"


c.

"WACC=0.55\times8(1-0.3)+0.45\times16=10.28"


"WACC=0.40\times8(1-0.3)+0.60\times16=11.84"